OpenStudy (anonymous):

Directions: write each equation in standard form using integers 36. y+2= 2/3(x+4)

OpenStudy (anonymous):

Soo standard form is y = mx+b do you know that?

OpenStudy (anonymous):

waaait no thats wrong, standard is when everything is on one side.

OpenStudy (anonymous):

y-2/3x = 32/12

OpenStudy (anonymous):

or 3y-2x = 8

OpenStudy (anonymous):

oh ok the formula is Ax+By=C

OpenStudy (anonymous):

Yep, solid. Good luck on the rest of them.

OpenStudy (anonymous):

oh ok which one woulds you go with